# Dissipation-Induced Steady States in Topological Superconductors: Mechanisms and Design Principles

### v1

@@ 5.4 The spectrum dependence on dissipation strength: the case $N_{M}=2$, $N_{B}=1$:
This behavior exemplifies non-Hermitian systems where perturbations create exceptional points and real-to-complex spectral transitions [114]. The parameter space near exceptional points is particularly noteworthy, though its analysis extends beyond this work.
@@ Acknowledgements
We are grateful to A. Abanov, E. König, Yu. Makhlin, and A. Mel’nikov for useful discussions and comments. The authors acknowledge the hospitality during the international workshop “Landau Week 2025” where part of this work has been performed.
@@
The authors acknowledge personal support from the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”. The work of I.S.B. was partially supported by the Basic Research Program of HSE. The work was supported by the Ministry of Science and Higher Education of the Russian Federation (Project No. FFWR-2024-0017).
7 Appendix
@@
$\!\!\!\!\hat{H}=\sum_{\nu=1}^{N}\varepsilon_{a}\hat{\alpha}_{a}^{+}\hat{\alpha}_{a},\qquad\hat{H}_{L}=\sum_{k}\xi_{k}\hat{d}_{k}^{+}\hat{d}_{k}$
@@
$\dot{\tilde{\rho}}_{T}(t)=-i\left[\,\tilde{V}(t)\,,\,\tilde{\rho}_{T}(t)\,\right],\qquad\tilde{\rho}_{T}(t)=\tilde{\rho}_{T}(-\infty)-i\int_{-\infty}^{t}ds\left[\,\tilde{V}(s),\leavevmode\nobreak\ \tilde{\rho}_{T}(s)\,\right].$ (83)
@@
$\tilde{c}_{n}(t)=\sum_{\nu=1}^{N}\left(\,u_{\nu n}^{*}\,e^{-i\varepsilon_{\nu}t}\,\hat{\alpha}_{\nu}+v_{\nu n}\,e^{i\varepsilon_{\nu}t}\,\hat{\alpha}^{+}_{\nu}\,\right)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{D}_{n}(t)=\sum_{k}t_{n,k}\,\hat{d}_{k}\,e^{-i\xi_{k}t}.$ (84)
@@
$\dot{\tilde{\rho}}(t)=-i\operatorname{Tr}_{E}\left(\left[\,\tilde{V}(t),\leavevmode\nobreak\ \hat{\rho}(-\infty)\otimes\,\hat{\rho}_{E}(-\infty)+\hat{\rho}_{corr}(-\infty)\right]\right)-\\ -\int_{-\infty}^{t}ds\operatorname{Tr}_{E}\left(\,\left[\,\tilde{V}(t),\,\left[\,\tilde{V}(s)\,,\leavevmode\nobreak\ \,\tilde{\rho}(t)\otimes\tilde{\rho}_{E}(t)+\tilde{\rho}_{corr}(t)\,\right]\,\right]\,\right)+\mathcal{O}(\,\hat{V}^{3}\,).$ (85)
@@
$\dot{\tilde{\rho}}(t)$ $=$ $\int_{0}^{\infty}ds\operatorname{Tr}_{E}\left(\,\left[\,\tilde{V}(t),\,\left[\,\tilde{V}(t-s)\,,\leavevmode\nobreak\ \,\tilde{\rho}(t)\otimes\hat{\rho}_{E}^{(eq)}\,\right]\,\right]\,\right)=$
$=$ $\int_{0}^{\infty}ds\,\sum_{n,m=1}^{N}\Big{(}\left[\,\tilde{c}_{n}(t)\,,\,\tilde{c}^{+}_{m}(t-s)\,\tilde{\rho}(t)\,\right]\cdot\operatorname{Tr}_{E}\left(\,\tilde{D}^{+}_{n}(t)\,\tilde{D}_{m}(t-s)\,\hat{\rho}_{E}^{(eq)}\right)+$
$\hphantom{aaaaaaaaal}+\,\left[\,\tilde{c}_{n}^{+}(t)\,,\,\tilde{c}_{m}(t-s)\,\tilde{\rho}(t)\,\right]\cdot\operatorname{Tr}_{E}\left(\,\tilde{D}_{n}(t)\,\tilde{D}^{+}_{m}(t-s)\,\hat{\rho}_{E}^{(eq)}\right)+h.c.\leavevmode\nobreak\ \,\Big{)}.$
The assumption (iii) implies that the correlation function of the reservoir involved in Eq. (7) depend on time $s$ only:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\operatorname{Tr}_{E}\left(\,\tilde{D}^{+}_{n}(t)\,\tilde{D}_{m}(t-s)\,\hat{\rho}_{E}^{(eq)}\right)=\operatorname{Tr}_{E}\left(\,\tilde{D}^{+}_{n}(s)\,\hat{D}_{m}\,\hat{\rho}_{E}^{(eq)}\right)=\sum_{k}t_{n,k}^{*}\,t_{m,k}\,e^{i\xi_{k}s}f(\xi_{k}),$ (87)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\operatorname{Tr}_{E}\left(\,\tilde{D}_{n}(t)\,\tilde{D}^{+}_{m}(t-s)\,\hat{\rho}_{E}^{(eq)}\right)=\operatorname{Tr}_{E}\left(\,\tilde{D}_{n}(s)\,\hat{D}^{+}_{m}\,\hat{\rho}_{E}^{(eq)}\right)=\sum_{k}t_{n,k}\,t_{m,k}^{*}\,e^{-i\xi_{k}s}f(-\xi_{k}).$ (88)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\dot{\hat{\rho}}(t)=-i\left[\,\hat{H}\,,\leavevmode\nobreak\ \hat{\rho}(t)\,\right]-\mathcal{D}[\,\hat{\rho}\,$]\,, (89)
@@
$A_{ml;\,a}=u_{am}u_{al}^{*}+v_{am}^{*}v_{al}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ B_{ml;\,a}=u_{am}v_{al}^{*}+v_{am}^{*}u_{al}\,$, (91)
@@
$\dot{\hat{\rho}}(t)$ $=$ $-i\left[\,\hat{H}+\hat{H}_{LS}\,,\leavevmode\nobreak\ \hat{\rho}(t)\,\right]+\mathcal{D}[\,\hat{\rho}\,$]\,, (98)
$\hat{H}_{LS}$ $=$ $\sum_{i,j=1}^{2N}\kappa_{ij}\,\hat{\Psi}_{i}\,\hat{\Psi}_{j}^{+}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{D}[\,\hat{\rho}\,]=\sum_{i,j=1}^{2N}\gamma_{ij}\left(\,2\,\hat{\Psi}_{i}^{+}\hat{\rho}\,\hat{\Psi}_{j}-\{\,\hat{\rho}\,,\,\hat{\Psi}_{i}\,\hat{\Psi}_{j}^{+}\,\}\,\right).$ (99)
@@
Here we assume that $\sqrt{\lambda_{v}}\in\mathbb{R}_{+}$ or $i\,\mathbb{R}_{+}$. Therefore, tunnel junction between the superconductor and the reservoir results in the GKSL equation with jump operators $\hat{L}_{v}$ which are linear in the fermionic operators. We note that the reservoir is equivalent to $\operatorname{rk}\gamma$ dissipative baths.

### v2

@@ 5.4 The spectrum dependence on dissipation strength: the case $N_{M}=2$, $N_{B}=1$:
This behavior exemplifies non-Hermitian systems where perturbations create exceptional points and real-to-complex spectral transitions [120]. The parameter space near exceptional points is particularly noteworthy and can lead to distinctive observable features [121], though its analysis extends beyond this work.
@@ Acknowledgements
We are grateful to A. Abanov, E. König, Yu. Makhlin, and A. Mel’nikov for useful discussions and comments. The authors acknowledge the hospitality during the international workshop ‘‘Landau Week 2025’’ where part of this work has been performed.
@@
The authors acknowledge personal support from the Foundation for the Advancement of Theoretical Physics and Mathematics ‘‘BASIS’’. The work of I.S.B. was partially supported by the Basic Research Program of HSE. The work was supported by the Ministry of Science and Higher Education of the Russian Federation (Project No. FFWR-2024-0017).
7 Appendices
7.1 Appendix A
@@
$\!\!\!\!\hat{H}=\sum_{{\color[rgb]{0,0,0}a=1}}^{N}\varepsilon_{a}\hat{\alpha}_{a}^{+}\hat{\alpha}_{a},\qquad\hat{H}_{L}=\sum_{k}\xi_{k}\hat{d}_{k}^{+}\hat{d}_{k}$
@@
$\dot{\tilde{\rho}}_{T}(t)=-i\left[\,\tilde{V}(t)\,,\,\tilde{\rho}_{T}(t)\,\right],\qquad\tilde{\rho}_{T}(t)=\tilde{\rho}_{T}(-\infty)-i\int_{-\infty}^{t}ds\left[\,\tilde{V}(s),\penalty 10000\ \tilde{\rho}_{T}(s)\,\right].$ (74)
@@
$\tilde{c}_{n}(t)=\sum_{\nu=1}^{N}\left(\,u_{\nu n}^{*}\,e^{-i\varepsilon_{\nu}t}\,\hat{\alpha}_{\nu}+v_{\nu n}\,e^{i\varepsilon_{\nu}t}\,\hat{\alpha}^{+}_{\nu}\,\right)\,,\penalty 10000\ \penalty 10000\ \tilde{D}_{n}(t)=\sum_{k}t_{n,k}\,\hat{d}_{k}\,e^{-i\xi_{k}t}.$ (75)
@@
[…]
@@
[…]
The assumption (iii) implies that the correlation function of the reservoir involved in Eq. (7.1) depend on time $s$ only:
[…]
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\dot{\hat{\rho}}(t)=-i\left[\,\hat{H}\,,\penalty 10000\ \hat{\rho}(t)\,\right]-\mathcal{D}[\,\hat{\rho}\,$]\,, (80)
@@
$A_{ml;\,a}=u_{am}u_{al}^{*}+v_{am}^{*}v_{al}\,,\penalty 10000\ \penalty 10000\ B_{ml;\,a}=u_{am}v_{al}^{*}+v_{am}^{*}u_{al}\,$, (82)
@@
$\dot{\hat{\rho}}(t)$ $=$ $-i\left[\,\hat{H}+\hat{H}_{LS}\,,\penalty 10000\ \hat{\rho}(t)\,\right]+\mathcal{D}[\,\hat{\rho}\,$]\,, (89)
$\hat{H}_{LS}$ $=$ $\sum_{i,j=1}^{2N}\kappa_{ij}\,\hat{\Psi}_{i}\,\hat{\Psi}_{j}^{+}\,,\penalty 10000\ \penalty 10000\ \mathcal{D}[\,\hat{\rho}\,]=\sum_{i,j=1}^{2N}\gamma_{ij}\left(\,2\,\hat{\Psi}_{i}^{+}\hat{\rho}\,\hat{\Psi}_{j}-\{\,\hat{\rho}\,,\,\hat{\Psi}_{i}\,\hat{\Psi}_{j}^{+}\,\}\,\right).$ (90)
@@
Here we assume that $\sqrt{\lambda_{v}}\geq 0$. It is important to note that since we did not employ the so-called rotating wave approximation, see, e.g., Ref.[39, 124], the matrix $\gamma$ is not guaranteed to be positive-definite. Therefore, the practical application of Eqs. (81)-(91) requires an additional analysis of positivity of $\sqrt{\lambda_{v}}$. Therefore, tunnel junction between the superconductor and the reservoir results in the GKSL equation with jump operators $\hat{L}_{v}$ which are linear in the fermionic operators. We note that the reservoir is equivalent to $\operatorname{rk}\gamma$ dissipative baths.
7.2 Appendix B
This Appendix presents the proof of Theorem 1 concerning the number and structure of zero kinetic modes in a system of free fermions under the influence of linear dissipative fields.
where the matrices $\mathcal{Z}_{0,1,2}$ read
we find for $\beta\to 0$:
$\mathcal{P}(\beta\to 0)\sim\beta^{2N_{M}-\operatorname{rk}\mathcal{Z}_{0}}\,,\,\penalty 10000\ \penalty 10000\ \penalty 10000\ \operatorname{rk}\mathcal{Z}_{0}\leq 2\min(N_{B},\,N_{M})\,.$ (100)
where we introduce
In the case $N_{M}>N_{B}$, the number of ZKM is bounded from below as
$N_{0}\geqslant 2(N_{M}-N_{B})$ (104)
The equality corresponds to the condition $\operatorname{rk}\mathcal{B}=2N_{B}$ and the following behavior of the characteristic polynomial $\mathcal{P}(\beta\to 0)\sim\beta^{2(N_{M}-N_{B})}$.
${\ell}\,\mathcal{B}^{T}\underline{y}_{a}=0,$ (107)
that is trivially satisfied due to $\underline{y}_{a}\in\ker\mathcal{B}^{T}$.
Since in Eq. (106) we construct eigen vectors corresponding to all existing ZKM, we can state that there is no other zero eigen vectors of $X$. Indeed, given that $\mathcal{B}\mathcal{B}^{T}\in\mathbb{R}^{2N_{M}\times 2N_{M}}$ and considering the result (103), we arrive at the relations:
$N_{0}=\dim\ker X=2N_{M}-\operatorname{rk}\mathcal{B}=\dim\ker\mathcal{\mathcal{B}\mathcal{B}}^{T}=\dim\ker\mathcal{B}^{T}.$
Thus, the dimension of the kernel of $\mathcal{B}^{T}$ coincides with the number of ZKMs, and therefore such solutions can only have the structure of the form (106).
7.3 Appendix C
The conditions of Theorem 1 imply the existence of $N_{M}$ strictly zero modes with $\varepsilon_{1}=\ldots=\varepsilon_{N_{M}}=0$ in the isolated system. These modes form a degenerate subspace, within which $N_{0}=2N_{M}-\operatorname{rk}\mathcal{B}$ zero kinetic modes with $\beta=0$ are realized in the dissipative system. In this Appendix, we examine the splitting of the ZKM subspace under a slight energy variation of the zero modes of isolated system.
We will assume that the spectrum of the isolated system contains $N_{M}$ subgap states with $0\leq\varepsilon_{b}\leq\delta$ ($b=1,\ldots,N_{M}$) and bulk states with $\Delta\leq\varepsilon_{c}\leq E$ ($c=N_{M}+1,\ldots,N$), where $\delta\ll\Delta$. In the case of $\delta=0$, the system’s characteristic polynomial $\mathcal{P}(\beta)=\det\left(X-\beta I\right)$ has a zero root of multiplicity $N_{0}$. Under small perturbations of the energy levels of the isolated system (of the order $\delta/\Delta\ll 1$), a correction $\sim\delta^{2}/\Delta^{2}$ arises in $\mathcal{P}(\beta)$, which leads to the splitting of the degenerate ZKM-subspace. Below, we estimate this splitting for a system with a single bath, $N_{B}=1$, as in this case all calculations can be performed explicitly.
It will be shown that for $N_{B}=1$, the characteristic polynomial $\mathcal{P}(\beta)$ behaves with respect to the small parameter $\delta/\Delta$ as follows (here we assume $N_{M}>1$; the case $N_{M}=1$ will be considered separately):
$\frac{1}{\Delta^{N}}\lim_{\,\delta/\Delta\to 0}\mathcal{P}(\beta)=\left(\frac{\beta}{\Delta}\right)^{2N_{M}-4}\left[\,\frac{\delta^{2}}{\Delta^{2}}\left(A+B\,\frac{\beta}{\Delta}\right)+\frac{\beta^{2}}{\Delta^{2}}R(\beta)\,\right]+O\left(\frac{\delta^{4}}{\Delta^{4}}\right),$ (108)
where the polynomial $R(\beta)$ of the power $2N-2N_{M}+2$ does not vanish at $\delta\to 0$, and $\lim_{\beta/\Delta\to 0}R(\beta)=\operatorname{const}$. It is easy to see that this asymptotic behavior corresponds to the polynomial
$\mathcal{P}\left(\beta\right)=\prod_{a=1}^{N_{M}-1}\left(\,\beta^{2}+\alpha_{a}^{2}\,\delta^{2}\,\right)\prod_{c=1}^{2(N-N_{M}+1)}\left(\,\beta-\alpha_{c}\,\Delta\,\right),\qquad\alpha_{a,c}\sim O(\delta^{0}).$ (109)
The form of this polynomial implies that $2N_{M}-2$ subgap modes have energies $\beta\sim i\delta$. Thus, the characteristic splitting of the ZKM energies is of the same order of magnitude as the splitting of the subgap isolated mode subspace, $|\,\beta\,|\sim\delta$.
To derive expression (108), we write the characteristic polynomial (94) in the form (where $d_{\beta}=\det\left(\,\mathcal{A}-\beta\,I\,\right)=\prod_{l=1}^{N}\left(\,\beta^{2}+\varepsilon_{l}^{2}\,\right)$ is the polynomial of the isolated system):
$\det(X-\beta\,I)=d_{\beta}^{-1}\cdot\det\left(\,I_{2}\cdot\prod_{l=1}^{N}\left(\,\beta^{2}+\varepsilon_{l}^{2}\,\right)-\sum_{l=1}^{N}\,\prod_{l^{\prime}\neq l}\left(\,\beta^{2}+\varepsilon_{l^{\prime}}^{2}\,\right)\left(\,\varepsilon_{l}\cdot Q^{(l)}+\beta\cdot P^{(l)}\,\right)\right),$ (110)
where matrices $Q^{(l)}$ and $P^{(l)}$ are defined in Eq. (94). Here,
$\det Q^{(l)}=\det P^{(l)}=\Gamma_{l}^{2}=\left(\,\big(\underline{\chi}_{2l-1}\cdot\underline{l}_{r}\big)\big(\underline{\chi}_{2l}\cdot\underline{l}_{i}\big)-\big(\underline{\chi}_{2l-1}\cdot\underline{l}_{i}\big)\big(\underline{\chi}_{2l}\cdot\underline{l}_{r}\big)\,\right)^{2}.$ (111)
The determinant on the right-hand side of (110) is divisible by the polynomial $d_{\beta}$ without remainder. Performing this division, we obtain the expression for the characteristic polynomial that determines the spectrum of kinetic modes:
[…]
Let us explicitly extract the coefficients $a_{p}$ of the powers $\beta^{p}$ using the following relations (higher-order derivatives can be computed using Faà di Bruno’s formula [125]):
[…]
In the last expression, the sum over $\{\,r\,\}_{j}$ means summation over all partitions of the number NjN-j into zeros and ones: $\{\,r\,\}_{j}=\{\,r_{1},\ldots,\,r_{N}=0,\,1:\penalty 10000\ r_{1}+\dots+r_{N}=N-j\,\}$  [^8] The product on the right-hand side of (113) can be visualized as a string (chain) with $N$ sites, where each site corresponds to a fermionic mode $\varepsilon_{l}$ of the closed system. This string contains $j$ vacancies, i.e., modes that are excluded from the product. Accordingly, the summation over $\{\,r\,\}_{j}$ corresponds to enumerating all possible arrangements of vacancies on a chain of length $N$. In cases where an additional $k$ elements are removed from the set, see Eq. (116), one should consider a chain of length NkN-k.. Then, we find the expression for the coefficients $a_{p}$ separately for even and odd powers of $\beta$:
$\begin{split}&a_{2j}=\sum_{\{\,r\,\}_{j}}\,\prod_{l=1}^{N}\varepsilon_{l}^{2r_{l}}+\frac{1}{2}\sum_{l_{1,2}=1}^{N}\left(\varepsilon_{l_{1}}\Gamma_{l_{2}}+\varepsilon_{l_{2}}\Gamma_{l_{1}}\right)^{2}\sum_{\{\,q\,\}^{\prime\prime}_{j}}\prod_{l^{\prime\prime}\neq l_{1,2}}\varepsilon_{l^{\prime\prime}}^{2q_{l^{\prime\prime}}}-\\ &\quad-(N-2)\sum_{l=1}^{N}\Gamma_{l}^{2}\,\sum_{\{\,s\,\}^{\prime}_{j}}\,\prod_{l^{\prime}\neq l}\varepsilon_{l^{\prime}}^{2s_{l^{\prime}}}+(2j^{2}-j)\det\left(P^{(l_{1})}+P^{(l_{2})}\right)\sum_{\{\,p\,\}^{\prime\prime}_{j-1}}\,\prod_{l^{\prime\prime}\neq l_{1,2}}\varepsilon_{l^{\prime\prime}}^{2p_{l^{\prime\prime}}},\end{split}$ (114)
[…]
Here, notations for different types of combinatorial distributions have been introduced:
$\begin{split}&\{\,r\,\}_{j}\penalty 10000\ \penalty 10000\ \,=\{\,r_{1},\ldots,\,r_{N}=0,\,1:\penalty 10000\ r_{1}+\dots+r_{N}=N-j\,\}\,;\\ &\{\,s\,\}^{\prime}_{j}\penalty 10000\ \penalty 10000\ \,=\{\,s_{1},\ldots,[s_{l}],\ldots,\,s_{N}=0,\,1:\penalty 10000\ s_{1}+\dots+s_{N}=N-j\,\}\,;\\ &\{\,p\,\}^{\prime\prime}_{j-1}=\{\,p_{1},\ldots,[p_{l_{1}}],\ldots,[p_{l_{2}}],\ldots\,p_{N}=0,\,1:\penalty 10000\ p_{1}+\dots+p_{N}=N-j+1\,\}\,;\\ &\{\,q\,\}^{\prime\prime}_{j}\penalty 10000\ \penalty 10000\ \,=\{\,q_{1},\ldots,[q_{l_{1}}],\ldots,[q_{l_{2}}],\ldots\,q_{N}=0,\,1:\penalty 10000\ q_{1}+\dots+q_{N}=N-j\,\}\,.\end{split}$ (116)
Variables in square brackets, for example $p_{l}\,]$, indicate that they are removed from the set. Thus, in products of the form $\prod_{l}\varepsilon_{l}^{2r_{l}}$ in Eqs. (114)-(115), not all energy modes are included as factors. When $N_{M}$ modes with $|\varepsilon|\sim\delta$ are present, this leads to different terms in the expressions for $a_{2j}$ having different asymptotics at small $\delta$. A straightforward analysis shows that in the generic cases where $\Gamma_{l}\neq 0$, the leading contribution to $a_{2j}$ is determined by the last terms in (114). Therefore,
$a_{2j-1}\sim\left(\frac{\delta}{\Delta}\right)^{\,\max\{2N_{M}-2j,\,0\}}\,,\qquad a_{2j}\sim\left(\frac{\delta}{\Delta}\right)^{\,\max\{2N_{M}-2j-2,\,0\}},\qquad\delta/\Delta\to 0.$ (117)
Then, assuming that $N_{M}\geq 2$ and normalizing all energy parameters by the gap $\Delta$, we obtain ( $\tilde{\beta}=\beta/\Delta,\quad b_{p}=a_{p}/\Delta^{N-p}$ ):
$\frac{1}{\Delta^{N}}\mathcal{P}(\beta)$ (118)
$=\underbrace{b_{0}+b_{1}\,\tilde{\beta}+\ldots}_{O\left(\delta^{4}/\Delta^{4}\right)}+\underbrace{b_{2N_{M}-4}\,\tilde{\beta}^{2N_{M}-4}+b_{2N_{M}-3}\,\tilde{\beta}^{2N_{M}-3}}_{O\left(\delta^{2}/\Delta^{2}\right)}+\underbrace{b_{2N_{M}-2}\,\tilde{\beta}^{2N_{M}-2}+\ldots+b_{2N}\,\tilde{\beta}^{2N}}_{O\left(\delta^{0}/\Delta^{0}\right)}.$
When analyzing the splitting features of the degenerate ZKM subspace at small $\delta$, the main contribution to $\mathcal{P}(\beta)$ will come from the terms in (118) of order $\sim O(\delta^{0}/\Delta^{0})$ and $\sim O\left(\delta^{2}/\Delta^{2}\right)$ As a result, we arrive at (108) and the conclusion that the splitting energies of the robust ZKMs are comparable to those of the MMs, $|\beta|\sim\delta$.
Figure 5: Dependence of subgap excitation energies on chain length $N$ for the isolated (hollow circles) and open (squares and dots) Kitaev chains (Sec. 5.1) with parameters $\mu=0.5$, $t_{r}=\delta_{r,N_{M}}$, $\Delta_{r}=0.8\,\delta_{r,N_{M}}$ (all energies are in units of $|t|$). (a,b) Homogeneous systems with $N_{M}=2$ and $N_{M}=3$, respectively. (c) Disordered system with $N_{M}=2$, where on-site energies $\epsilon_{j}$ (Eq. (59)) are drawn from a Gaussian distribution (mean $\bar{\epsilon}=0$ and dispersion $\sigma_{\epsilon}=\mu/2$). The energies $\beta$ for the open system are averaged over 500 random realizations of the dissipation fields $\underline{\mu},\underline{\nu}\in\mathbb{C}^{N}$.
We verified these predictions numerically in the framework of the extended Kitaev chain model from Sec. 5. Figure 5 shows the dependence of subgap excitation energies on the chain length $N$ for both the isolated ($0<\varepsilon\ll\Delta$, hollow circles) and open ($|\beta|\ll\Delta$, squares/dots) systems. Panels (a,b) display homogeneous systems with topological indices $N_{M}=2,3$999 Recall that $N_{M}$ is the number of zero-mode pairs in the thermodynamic limit.. Panel (c) shows a disordered case with $N_{M}=2$ and random on-site energies $\epsilon_{j}$ (Eq. (59)) drawn from a Gaussian distribution ($\bar{\epsilon}=0$, $\sigma_{\epsilon}=\mu/2$). The open system energies $\beta$ were averaged over 500 random realizations of dissipation fields $\underline{\mu},\underline{\nu}\in\mathbb{C}^{N}$.
Figure 5 exhibits that finite-size effects cause Majorana mode wave functions to overlap, generating the well-known subgap excitations with exponentially decaying energy $\varepsilon\sim e^{-N}$ [114]. In homogeneous systems, when $N\,/\,N_{M}\in\mathbb{N}$, all subgap modes acquire the same energy $\varepsilon_{1}=\ldots=\varepsilon_{N_{M}}$ (panels a,b). This degeneracy is lifted in the inhomogeneous case ($\varepsilon_{1}\neq\varepsilon_{2}$ on panel c). The dependencies demonstrate that the subgap mode energies of the isolated and open systems are comparable. Furthermore, it is clear that this conclusion holds in the presence of disorder originating from both the subsystem itself and the dissipative fields.
While the above analysis pertained to robust ZKMs, we now turn to the features of weak ZKMs, considering the simplest possible case $N_{M}=N_{B}=1$. In this case all coefficients of the polynomial $\mathcal{P}(\beta)$ have a finite limit as $\delta/\Delta\to 0$. Accordingly, robust ZKMs are not realized, which is consistent with Theorem 1.
However, in addition to the robust zeros, weak ones can arise due to the hybridization of the closed system’s wave functions with dissipative fields. To demonstrate this, consider the expression for the constant term of the polynomial $a_{0}$:
$a_{0}=\prod_{l=1}^{N}\varepsilon_{l}^{2}+\left[\,\sum_{l=1}^{N}\left(\,\big(\underline{\chi}_{2l-1}\cdot\underline{l}_{r}\big)\big(\underline{\chi}_{2l}\cdot\underline{l}_{i}\big)-\big(\underline{\chi}_{2l-1}\cdot\underline{l}_{i}\big)\big(\underline{\chi}_{2l}\cdot\underline{l}_{r}\big)\,\right)\cdot\prod_{l^{\prime}\neq l}\varepsilon_{l^{\prime}}\,\right]^{2}.$ (119)
The equality $a_{0}=0$ means that $\beta=0$ is a solution of the equation $\det(X-\beta I)$ with multiplicity one. Next, we assume that $\varepsilon_{1}=0$. Then, the existence of a weak zero is determined by the condition
$a_{0}\sim\Gamma_{1}=\big(\underline{\chi}_{1}\cdot\underline{l}_{r}\big)\big(\underline{\chi}_{2}\cdot\underline{l}_{i}\big)-\big(\underline{\chi}_{1}\cdot\underline{l}_{i}\big)\big(\underline{\chi}_{2}\cdot\underline{l}_{r}\big)=0,$ (120)
which agrees with Eq. (66). Next, maintaining the conditions for the realization of a weak ZKM for the case $\varepsilon_{1}=0$ (assuming $\Gamma_{1}=0$), let us suppose that the spectrum of the isolated system is perturbed such that $\varepsilon_{1}\sim\delta$. Then, from the analysis of (120), it follows that $a_{0}\sim\delta^{2}/\Delta^{2}$, and hence the splitting of the weak ZKM behaves as $|\,\beta\,|\sim\delta^{2}/\Delta$ for $\delta/\Delta\to 0$. Thus, in dissipative systems with a single bath $N_{B}=1$, weak zeros are more stable to small perturbations of the isolated system’s spectrum than robust ZKMs.
The presented analysis allows us to assert the robustness of the ZKMs described in Sec. 4 against small local perturbations of the Hamiltonian or dissipative fields in the case of $N_{B}=1$, and to suggest a similar robustness for cases with $N_{B}>1$.